If the volume of the box is 1600 cubic centimeters, what are the dimensions of the box if no side can be greater than 25 centimeters?
Answer by richwmiller(17219)

vol=l*w*h
What are the factors of 1600
100*16
We need three factors with no factor > 25
100*16
25*4*(2*2*4) =25*4*16
or
25*(4*2)*(4*2)=25*8*8